OpenStudy (anonymous):

Convert the mass of water to moles of water 0.529 g PLEASE EXPLAIN

OpenStudy (kainui):

Alright so I'll help break it down nice and easy, because this is what the numbers on the periodic table are about. There are two numbers, right? The smaller one is the number of protons, and that defines what the atom IS. So when you see 8 for Oxygen, that means all oxygen atoms have 8 protons no matter what. If it had 7 it would be called Nitrogen and if it had 9 it would be called Fluorine. We didn't know that all along we were calling small clusters of protons names. So this is a little long winded, but bear with me, cause it's important and slightly interesting if you want to understand reality! Alright so now the atom's nucleus has neutrons there as well, but they don't matter what number they are quite as much. Different numbers of neutrons are called isotopes of the same element and atoms usually prefer to have a certain number of these guys hanging around. Usually they are about as many neutrons as protons and they weigh the same. So when you see the other number, 15.999 near oxygen, that's because there's 16 total nucleons (nucleons are what neutrons and protons are called because they're the same size and in the same place unlike electrons which are smaller and "floating" around, nothing fancy.) BUT that's not the end of the story. See, since it's 16 nucleons that means you have 8 protons and 8 neutrons right? Well that number isn't 16, it's 15.999! That's because it's the average weight of a mole of that type of atom. So some of the oxygen in a big grab-bag of oxygen will have 7 neutrons instead of 8, but it's very rare. This is what you'll be using since you have just normal water, some of the atoms are made up of these different isotopes of oxygen and hydrogen. That means if you have 15.999 grams of Oxygen atoms, it means there is 1 mole of them in the pile. \[\frac{ 15.999 g }{ 1mol }O atoms\] This is basically your conversion factor telling you how many grams per mole. Just turn the fraction on its head to see how many moles per gram you have, and make sure you cancel out the units correctly. You had water though, not oxygen, so you need to make sure you add up the right amounts, which is 1 mole of oxygen and 2 moles of hydrogen for every 1 mole of water. 15.999 g/mol O atoms+2*1.008 g/mol H atoms=18.015 grams per mole of water. Remember that to make water you need to go through a reaction kind of like this: \[O+2H \rightarrow H_2O\] (chemistry nerds, don't hate on this, not going to make this excessively complicated) Then make sure the units cancel out to 1: \[0.529 g H_2O * \frac{ 1 mol H_2O }{ 18.015 gH_2O }=molH_2O\] I can help clarify anything if you need it... And hopefully you don't hate me haha.